# Rethinking the Integration of Prediction and Planning in Deep Learning-Based Automated Driving Systems: A Review

### v2

@@ Abstract
Automated driving has the potential to revolutionize personal, public, and freight mobility. Beside accurately perceiving the environment, automated vehicles must plan a safe, comfortable, and efficient motion trajectory. To promote safety and progress, many works rely on modules that predict the future motion of surrounding traffic. Modular automated driving systems commonly handle prediction and planning as sequential, separate tasks. While this accounts for the influence of surrounding traffic on the ego vehicle, it fails to anticipate the reactions of traffic participants to the ego vehicle’s behavior. Recent models increasingly integrate prediction and planning in a joint or interdependent step to model bi-directional interactions. To date, a comprehensive overview of different integration principles is lacking. We systematically review state-of-the-art deep learning-based prediction and planning, and focus on integrated prediction and planning models. Different facets of the integration ranging from model architecture and model design to behavioral aspects are considered and related to each other. Moreover, we discuss the implications, strengths, and limitations of different integration principles. By pointing out research gaps, describing relevant future challenges, and highlighting trends in the research field, we identify promising directions for future research.
@@
6.6in(0.95in,0.15in) SUBMITTED TO REVIEW AND POSSIBLE PUBLICATION. COPYRIGHT WILL BE TRANSFERRED WITHOUT NOTICE.
@@ 1 Introduction
Automated Driving (AD) remains a challenging endeavour. It is usually split into the subtasks of perception, prediction, planning, and control [1, 2, 3]. Perception processes sensor inputs to create a model of the environment. Prediction and planning build upon this model and make future motion forecasts for surrounding traffic agents and a plan for the controlled ego vehicle. Traditional, modular systems (cf. Fig. 1) address prediction and planning as separate tasks. The predicted behavior of surrounding traffic agents is leveraged to plan a suitable behavior for the ego vehicle. However, this sequential ordering is inherently reactive and cannot represent bidirectional interaction between the ego vehicle and other traffic agents [4]. In fact, prediction and planning are no sequential problems and should be tightly coupled in automated driving systems [5, 6, 7].
1.1 Scope
In this work, we review methods to integrate prediction and planning into an automated driving system (ADS).
Prediction is the task of anticipating intents and future trajectories of observed traffic agents [8], and planning is about finding the best possible trajectory w.r.t. previously defined criteria for a controlled vehicle [9].
@@ 1.1 Scope
Deep Learning (DL)-based methods follow a data-centric approach to tackle these problems [10] and have led to significant improvements in many fields [11]. Works in prediction and planning increasingly employ DL-based approaches as well [12, 13, 14, 15] and represent an alternative to traditional, rule-based methods [16, 17] and constrained optimization [18]. In this work, we focus on DL-based methods.
@@
The architecture of an ADS can take different forms, which are shown in Fig 1. Black-box methods employ a single neural network for perception, prediction and planning (c.f. Fig 1(a)), so that depending on the architecture, no boundaries between the modules can be drawn [9, 19, 20]. In contrast, modular systems consist of clearly defined components for different tasks [21, 22, 23, 24, 25]. Traditionally, they are arranged in a sequential order (c.f. Fig 1(b)). In this work, we emphasize the shortcomings of this approach and highlight works that integrate prediction and planning in a more sophisticated way (c.f. Fig 1(c)) to enable bi-directional interaction of self-driving vehicles and surrounding traffic. We refer to the component of an automated driving system that integrates prediction and planning as an Integrated Prediction and Planning System (IPPS).
We limit the scope of our work to mixed traffic without Vehicle-to-X communication between traffic agents. Furthermore, we exclude pedestrian motion forecasting, which poses a different problem that is subject to weaker dynamic constraints compared to vehicles [26] and has been thoroughly surveyed [27, 28, 29].
1.2 Contribution and Structure
In a thorough approach, this survey first reviews DL-based prediction and planning before focusing on how to integrate these two tasks. While surveys on classical methods [30, 31, 14, 32], stand-alone DL-based prediction [33, 34, 35, 15] or planning [36], and end-to-end (E2E) AD [37] already exist, to the best of our knowledge we are the first to review the integration of DL-based prediction and planning. We summarize our contributions as follows:
  • We provide a comprehensive overview of common representations, architectures and state-of-the-art methods for DL-based prediction (Sec. 3) and planning (Sec. 4).
@@ 1.2 Contribution and Structure
  • We propose a categorization for the integration of prediction and planning, which addresses model architecture, system design and behavioral aspects. Further, we investigate and analyze the interrelations between these categories and their impact on safety and robustness (Sec. 5).
@@
  • We reveal gaps in state-of-the-art research and point out promising directions of future research based on the identified categorization (Sec. 6).
(a) Black box ADS
(b) ADS with sequential IPPS
(c) ADS with bidirectional IPPS
Figure 1: System architectures of automated driving systems. Dashed lines indicate, that due to the black-box nature of some E2E ADS, it can be unclear which network components do prediction and planning, so that no boundary around the IPPS can be drawn.
2 Task Definitions
@@
and the future states of all SVs are $\overline{\textbf{Y}}_{\textrm{SV}}$. In the following, state sequences are also referred to as trajectories.
Additional scene information, such as a semantic map or traffic signs and traffic light states are represented by $I$.
Prediction. Following [8], we state trajectory forecasting as the task of estimating a probability distribution
@@
that maps state histories of $m$ observed vehicles in $\overline{\textbf{X}}$ to the future trajectories of $n$ predicted vehicles in $\overline{\textbf{Y}}$. The distribution $P_{\textrm{pred}}$ accounts for inherent uncertainties in the forecasting task and is often modeled by a discrete set of samples with corresponding probabilities [38, 39, 40, 41]. Some methods omit scene information $I$ and infer trajectories based on a state history alone [42, 43, 44].
TABLE I: Forms of Prediction
$\overline{\textbf{Y}}$ $\overline{\textbf{X}}$ Task
$Y_{\textrm{SV}}$ $\overline{\textbf{X}}_{\textrm{SV}}$, $X_{\textrm{EV}}$ Single-Agent Prediction
$\overline{\textbf{Y}}_{\textrm{CV}}$ $\overline{\textbf{X}}_{\textrm{SV}}$, $X_{\textrm{EV}}$ Clique Prediction
$\overline{\textbf{Y}}_{\textrm{SV}}$ $\overline{\textbf{X}}_{\textrm{SV}}$, $X_{\textrm{EV}}$ Joint Prediction
[…]
Planning. The task is to find a single suitable trajectory for the EV that can be passed on to a downstream motion controller. Hence, we define planning to be a function $f$ that maps the observational inputs $\overline{\textbf{X}}_{\textrm{SV}}$, and $X_{\textrm{EV}}$ as well as the context information $I$ to a future trajectory $Y_{\textrm{EV}}$, i.e.
@@
The definitions show that planning can be considered a special case of single-agent prediction where the output distribution models only a single trajectory. However, in contrast to prediction, the planning trajectory must be conditioned on a navigation goal. Moreover, it has to be stable in closed-loop, i.e. it must be kinematically feasible and result in safe and efficient driving behavior when fed to a downstream controller.
3 Prediction
The core of prediction is understanding how the driving scene will evolve. Sec. 3.1 discusses how different scene representations allow local and global interactions. Sec. 3.2 reviews which NN designs are used to model interactions and extract descriptive features. Finally, Sec. 3.3 shows how the extracted features are mapped to trajectory predictions and how multimodality is modeled.
3.1 Scene Representation
Figure 2: Comparison of rasterized and sparse representation of the same traffic scene. While the former represents information with fixed-size grids, the latter uses a set of vectors.
@@ 3.1 Scene Representation
Representing a scene means extracting the relevant subset of all available information and transferring it into a format that subsequent processing steps can utilize. In the context of trajectory prediction for automated vehicles, agent states $\overline{\textbf{X}}$ and map I are the most important information to represent. Two major representations exist in DL-based approaches: Rasterized and sparse (cf. Fig. 2).
Rasterized. Rasterized representations use dense, fixed-resolution grid structures that often have multiple channels [49, 50, 48, 51, 52]. Each channel encodes different information on agent states. They are commonly used in combination with rasterized high-definition maps (HD maps). Typical modalities are occupancy maps [51], object detections [53], lane markings [54], motion history [55], semantic information [56], traffic signs [57], traffic light states [57], and route information [58] (c.f. Fig 2.
DESIRE [8] is an early work that uses a rasterized representation for DL-based prediction. A bird’s eye view (BEV) representation is employed to fuse multiple EV sensor inputs into a shared representation. This allows the combination of different sensing modalities and establishes a common coordinate system for all vehicles, facilitating to model interactions within the scene [59].
Additionally, dense grids are well-suited for processing with powerful Convolutional Neural Network (CNN) [60] architectures [58, 61, 62, 10, 63, 64, 65, 66, 67, 68, 69, 70, 71]. However, embedding all observations in a grid structure leads to information loss due to quantization errors [72]. Moreover, as discussed in Sec. 3.2, the locally restricted receptive field of CNNs can hamper interaction modeling.
@@
Sparse representations use vectors to describe all objects in a scene. The per-object vectors are then either encoded jointly, e.g., with a transformer, or individually and then aggregated, e.g., with Graph Neural Networks (GNN) [73]. Objects are represented by polygons, or point sets [6, 74, 41, 75, 61, 76, 38, 77]. The map is commonly described by a set of lanes represented by polylines and an adjacency matrix. Objects and map elements are then encoded into fixed-size latent features, for instance, via Multi-Layer Perceptrons (MLP) [78] or Recurrent Neural Networks (RNN) [79, 39]. To encode sequential signals like state histories $\overline{\textbf{X}}$, 1D CNNs [80], RNNs like LSTMs [81] and GRUs [82], and Transformers [83] are applied [41, 84, 76].
The timeline in Fig. 3 chronologically depicts impactful prediction methods and exhibits a shift towards sparse representations during the past years. Diehl et al. [72] proposed a graph representation for vehicle interaction modeling. Based on their work, two approaches emerged: Homogeneous and heterogeneous graphs. Homogeneous graphs encode objects of one class, for example, the road network [74, 41], whereas heterogeneous graphs combine different object classes in one graph [40, 85, 86, 87]. However, their work is focused on highway data and does not include map information. In contrast, VectorNet [74] proposes to represent all map elements and agents by a set of points. Agents are represented by a set of points along their motion history, while lanes are described by polyline points. Small objects, such as traffic signs, are represented by a single point, and extended elements are given by polygons defined by multiple points. Vectornet then processes each object with a homogeneous subgraph called polyline subgraph. A per-object feature is obtained by max-pooling. Next, all object features are aggregated in a heterogeneous global graph. Finally, the TV trajectory is decoded from the encoding of the node corresponding to the TV. Trajectron++ [40] and MFP [88] combine rasterized representations for HD maps with sparse representations for objects. Further, a few works use voxel representations that differ in their degree of sparsity [89][90].
MultiPath++ [38] is the successor of MultiPath [56] and switches from rasterized to sparse representations. Their results show that the shift from rasterized to sparse can increase performance. The overall development in the field confirms this trend (c.f. Fig. 3).
Coordinate frame. Irrespective of the scene representation itself, the coordinate frame is a fundamental design choice.
Applying a global coordinate system with a fixed viewpoint for the whole scene [91, 46, 92, 6] is computationally efficient. However, the frame is not viewpoint-invariant, i.e., the predictions change if the coordinate system is moved within the scene. This leads to low sample efficiency and impairs generalization [93]. Even fixing the coordinate origin to the EV center does not lead to viewpoint-invariant predictions when predicting other agents than the EV [94].
In contrast, other models [95, 96, 74, 54] process a scene from the viewpoint of each agent. Thus, the processing is viewpoint-invariant. On the downside, computational complexity scales linearly with the number of agents and quadratically with the number of interactions [6].

### v3

# The Integration of Prediction and Planning in Deep Learning Automated Driving Systems: A Review

@@ Abstract
Automated driving has the potential to revolutionize personal, public, and freight mobility. Beside accurately perceiving the environment, automated vehicles must plan a safe, comfortable, and efficient motion trajectory. To promote safety and progress, many works rely on modules that predict the future motion of surrounding traffic. Modular automated driving systems commonly handle prediction and planning as sequential, separate tasks. While this accounts for the influence of surrounding traffic on the ego vehicle, it fails to anticipate the reactions of traffic participants to the ego vehicle’s behavior. Recent methods increasingly integrate prediction and planning in a joint or interdependent step to model bidirectional interactions. To date, a comprehensive overview of different integration principles is lacking. We systematically review state-of-the-art deep learning-based planning systems, and focus on how they integrate prediction. Different facets of the integration ranging from system architecture to high-level behavioral aspects are considered and related to each other. Moreover, we discuss the implications, strengths, and limitations of different integration principles. By pointing out research gaps, describing relevant future challenges, and highlighting trends in the research field, we identify promising directions for future research.
@@
Note: A beginner-friendly introduction to prediction, planning, and their integration in automated driving systems can be found in an earlier version of our work at https://arxiv.org/abs/2308.05731v2. The following version is intended for more experienced researchers in the field.
@@
7in(0.75in,0.15in) (c) IEEE 2024.
@@ I Introduction
Automated Driving (AD) remains a challenging endeavour. It is usually split into the subtasks of perception, prediction, planning, and control [1, 2, 3]. Perception processes sensor inputs to create a model of the environment. Prediction and planning build upon this model and make future motion forecasts for surrounding traffic agents and a plan for the controlled ego vehicle. Traditional, modular systems (cf. Fig. 2) address prediction and planning as separate tasks. The predicted behavior of surrounding traffic agents is leveraged to plan a suitable behavior for the ego vehicle. However, this sequential ordering is inherently reactive and cannot represent bidirectional interaction between the ego vehicle and other traffic agents [4]. In fact, prediction and planning are no sequential problems and should be tightly coupled in automated driving systems [5, 6, 7]. Fig. 1 highlights the interdependence of both tasks.
Figure 1: How should the ego vehicle behave? Various behaviors are possible for the ego vehicle. Here, we depict two in red and orange. Their consequences depend on the behavior of the observed surrounding vehicle, which is depicted by blue arrows (prediction A or prediction B). At the same time, the surrounding vehicle might react to the ego vehicle’s action, i.e. the surrounding vehicle’s behavior also depends on the ego vehicle’s decision. Different methods exist to forecast the behavior of the surrounding vehicle, and various ways exist to leverage this to decide on a safe and goal-directed plan for the ego-vehicle. In this survey, we systematically categorize and review methods that integrate prediction and planning for self-driving vehicles. We highlight their capabilities and limitations and discuss prospects for future research.
I-A Scope
@@
In this work, we review methods to integrate prediction and planning into an automated driving system. Prediction is the task of anticipating intents and future trajectories of observed traffic agents [8], and planning is about finding the best possible trajectory w.r.t. previously defined criteria for a controlled vehicle [9]. Deep Learning (DL)-based methods follow a data-centric approach to tackle these problems [10] and have led to significant improvements in many fields [11]. Works in prediction and planning increasingly employ DL-based approaches as well [12, 13, 14, 15] and represent an alternative to traditional, rule-based methods [16, 17, 18] and constrained optimization [19]. In this work, we focus on DL-based methods.
@@
An overview of different architectures of automated driving systems is shown in Fig. 2. It can take three forms ranging from modular systems over end-to-end (E2E) differentiable modular systems to monolithic E2E systems. Modular and modular E2E systems consist of clearly defined components for different tasks [20, 21, 22, 23, 24]. Traditionally, they are arranged sequentially. We will discuss these systems in Sec. III. In contrast, monolithic E2E methods employ a single neural network for perception, prediction, and planning so that depending on the architecture, no boundaries between the modules can be drawn [9, 25, 26]. As the order of the subtasks cannot be determined, we name this integration of prediction and planning “undirected” (c.f. Sec. IV). In this work, we emphasize the shortcomings of sequential and undirected approaches and highlight works that integrate prediction and planning in a more sophisticated way to enable bidirectional interaction of self-driving vehicles and surrounding traffic (c.f. Sec. V). In the following, we refer to the component of an automated driving system that integrates prediction and planning as an Integrated Prediction and Planning System (IPPS).
Figure 2: Overview of automated driving systems. There are three system architectures shown on the left. Modular systems consist of individual modules, whose interfaces provide interpretability but restrict information flow and end-to-end differentiability. In contrast, monolithic E2E systems are end-to-end differentiable but not interpretable. Modular E2E systems combine both properties: they are end-to-end differentiable and interpretable. All three system architectures comprise an integrated prediction and planning system (IPPS), depicted in green. Our work focuses on this very part and identifies three paradigms to integrate prediction and planning, as shown on the right. Sequential IPPSs condition one task on the other. Undirected IPPSs allow for more complex interactions but provide low interpretability. Bidirectional IPPSs explicitly ensure that both tasks are mutually conditioned on each other. Our review analyzes these integration paradigms and highlights their compatibility with different system architectures.
I-B Contribution and Outline
By discussing automated driving systems with a focus on the integration of prediction and planning, we take a novel perspective. IPPSs can be classified into three categories as depicted in Fig. 2. Accordingly, this survey is structured into three main chapters: sequential (Sec. III), undirected (Sec. IV), and bidirectional IPPSs (Sec. V). While surveys on classical methods [27, 28, 14, 29], stand-alone DL-based prediction [30, 31, 32, 15] or planning [33], and E2E AD [34] already exist, we observe that recent methods thoughtfully design the interplay of prediction and planning [35, 36, 37, 38]. To substantiate this observation with broad evidence and a theoretical foundation, we summarize our contributions as follows:
@@ I-B Contribution and Outline
  • We propose a categorization for the integration of prediction and planning based on the dependencies between both tasks (cf. Sec. IIIIV and V). We investigate how these categories relate to system architectures, behavioral aspects, and safety.
  • In particular, we provide a comprehensive overview of design choices in prediction and planning modules of modular (E2E) architectures (cf. Sec. III-A and III-B) and discuss their impact on interaction and system-level behavior.
@@
  • We reveal gaps in state-of-the-art research and point out promising directions of future research based on the identified categorization (Sec. VI).
II Task Definitions
@@
and the future states of all SVs are $\overline{\textbf{Y}}_{\textrm{SV}}$. In the following, state sequences are also referred to as trajectories. Additional scene information, such as a semantic map or traffic signs and traffic light states are represented by $I$.
Following [8], we state trajectory prediction as the task of estimating a probability distribution
@@
that maps state histories of $m$ observed vehicles in $\overline{\textbf{X}}$ to the future trajectories of $n$ predicted vehicles in $\overline{\textbf{Y}}$. The distribution $P_{\textrm{pred}}$ accounts for inherent uncertainties in the forecasting task and is often modeled by a discrete set of samples with corresponding probabilities [39, 40, 41, 42]. Some methods omit scene information $I$ and infer trajectories based on a state history alone [43, 44, 45]. Commonly $\overline{\textbf{X}}$ includes all vehicles in the scene, i.e. $X_{\textrm{EV}}$ and $\overline{\textbf{X}}_{\textrm{SV}}$. Depending on the vehicles in $\overline{\textbf{Y}}$, different variants of prediction can be formulated: single-agent prediction models the future trajectory for each SV individually [46]. However, estimating a joint prediction $\overline{\textbf{Y}}_{\textrm{SV}}$ from a set of $n$ single-agent predictions $Y_{\textrm{SV}}$ is not trivial since the number of possible combinations grows exponentially with the number of agents $n$. Not all combinations are meaningful, and finding realistic ones with heuristics and joint optimization is cumbersome [47]. To avoid this, joint prediction directly estimates a joint distribution for multiple SVs [48].
In planning, the task is to find a single suitable trajectory for the EV that can be passed on to a downstream motion controller. Hence, we define planning to be a function $f$ that maps the observational inputs $\overline{\textbf{X}}_{\textrm{SV}}$, and $X_{\textrm{EV}}$ as well as the context information $I$ to a future trajectory $Y_{\textrm{EV}}$, i.e.
@@
The definitions show that planning can be considered a special case of single-agent prediction where the output distribution models only a single trajectory. However, in contrast to prediction, the planning trajectory must be conditioned on a navigation goal. Moreover, it has to be stable in closed-loop, i.e. it must be kinematically feasible and result in safe and efficient driving behavior when fed to a downstream controller [49].
III Sequential IPPS
To date, the sequential integration of prediction and planning is the prevalent and most widely adopted integration principle. Sequential IPPSs execute prediction and planning as separate tasks. Prior knowledge is used to design the interplay of the modules. Depending on the interfaces, such systems can still be end-to-end differentiable. Accordingly, sequential integration can be found in modular and modular E2E systems (c.f. Fig. 2). To investigate different sequential system designs in more depth, we start by taking a closer look at the individual components of prediction and planning and discuss their impact on interaction and system-level behavior.
III-A Prediction
III-A1 Input Representations
@@
In the context of IPPSs for automated vehicles, agent states $\overline{\textbf{X}}$ and map I are the most important information to represent. Two major representations exist in DL-based approaches: Rasterized and sparse.
Rasterized representations use dense, fixed-resolution grid structures that often have multiple channels [50, 51, 52, 53, 54, 55]. Each channel encodes different information on agent states. They are commonly used in combination with rasterized high-definition maps (HD maps) [56, 57] and represented in a bird’s eye view (BEV) [58, 59]. The BEV allows to fuse multiple EV sensor inputs and different sensing modalities into a shared and interpretable representation [8]. It further establishes a common coordinate system for all vehicles, facilitating to model interactions within the scene [60]. Dense grids are well-suited for processing with powerful Convolutional Neural Network (CNN) [61] architectures [56, 62, 63, 10, 64, 65, 66, 67, 68, 69, 70, 71, 72]. However, embedding all observations in a grid structure leads to information loss due to quantization errors [73]. Moreover, the locally restricted receptive field and the limited resolution of CNNs can hamper interaction modeling.
@@
Sparse representations use vectors to describe all objects in a scene. The per-object vectors are then either encoded into a latent representation jointly, e.g., with a Transformer, or individually and then aggregated, e.g., with Graph Neural Networks (GNN) [74]. Objects are represented by polygons, or point sets [6, 75, 42, 76, 62, 77, 39, 78, 79, 80, 40, 81, 82]. The map is commonly described by a set of lanes represented by polylines and an adjacency matrix. Objects and map elements are then encoded into fixed-size latent features, for instance, via Multi-Layer Perceptrons (MLP) [83] or Recurrent Neural Networks (RNN) [84]. To encode sequential signals like state histories $\overline{\textbf{X}}$, 1D CNNs [85], RNNs like LSTMs [86] and GRUs [87], and Transformers [88] are applied [42, 89, 77]. Some works like Trajectron++ [41] and MFP [90] combine rasterized HD maps with sparse object representations.
Throughout the past years, a clear trend towards sparse input representations can be observed in prediction modules [91, 92, 93, 94, 95, 96, 97, 98]. In the context of integrating prediction and planning, an advantage of sparse representations is that they are object-based. Interactions between the EV and SVs can therefore be modeled in a targeted manner. This improves the explainability of a system and provides a clear understanding of how surrounding vehicles influence the ego-plan.
III-A2 Interaction Modelling
Provided with an accurate scene representation, interaction modeling between perceived objects is a decisive factor for the successful integration of prediction and planning [4, 99]. Modeling interactions between vehicles helps to understand how the actions of a vehicle affect the behavior of other traffic participants [98]. This model is subject to various influences, including traffic rules, physics, and common sense reasoning. Interaction modeling also includes interactions between map elements and vehicles. For example, modeling the relation of vehicles to static map elements like lane markings, traffic signs, etc. is important to understand feasible corridors in which trajectories could be localized. Some models employ rule-based heuristics to guide interaction modeling, e.g., ScePT [52] only models interaction among nearby vehicles. Similarly, heterogeneous graph attention network (HGAT) [100] and graph-based interaction-aware trajectory prediction (GRIP) [101] use heuristics to identify which pairs of nodes in the scene graph are to be connected with edges. Other methods rely entirely on neural components to learn these relations. Altogether, an interaction model should yield an understanding of how the scene can potentially evolve in the future.
Recapitulating recent methods exposes that diverse architectures are used to realize interaction modeling. RNNs are only used in early prediction models like DESIRE [8] or Trajectron++ [41] and combined with an aggregation operator like spatial pooling or attention. Alternatively, CNN methods apply 2D convolutions to implicitly capture interaction within the kernel size [102, 103, 104]. Compared to sequence processing-based approaches, higher importance is assigned to spatial interaction. However, no explicit interactions between designated objects are modeled. This shortcoming is addressed by GNNs and Graph-Attention, which explicitly model interactions between individual agents. They either combine multiple agents’ features [105] and process them with graph convolution operators [106, 74, 101, 107] or perform graph attention [108, 109, 110, 111, 112] to aggregate information. Recently, Transformers [88] are widely adopted for interaction modeling due to their global receptive field and attention mechanism [113, 77, 114, 6, 91, 92, 93, 95, 97, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124]. All SVs can be predicted simultaneously [77], and the impact of vehicles on each others’ behaviors across different timesteps can be modeled by masking [114]. Attention is applicable along the spatial and temporal dimension jointly [114], or individually [125]. The latter break down into sequential [126, 127] and interleaved [6, 92] attention. Experiments with Wayformer [92] and SceneTransformer [6] show that joint and interleaved processing outperform sequential approaches. As joint processing causes a higher computational burden, the interleaved paradigm balances best between performance and computational demand. This property of Transformers can provide explainability to the integration of prediction and planning. Attention can be utilized flexibly to achieve locality restrictions, causality, and an explicit interaction design. For example, a sequential IPPS could first model interactions between SVs and then include the EV for planning.
III-A3 Coordinate Frame
Irrespective of the scene representation itself, the coordinate frame is a fundamental design choice that affects interaction modeling and thus the integration of prediction and planning. Applying a global coordinate system with a fixed viewpoint for the whole scene [128, 47, 129, 6] is computationally efficient as it can be shared across both modules. However, the frame is not viewpoint-invariant, i.e., the predictions and the plan change if the origin of the coordinate system is moved within the scene. This leads to low sample efficiency and impairs generalization [130]. Therefore, the origin is commonly fixed to the center of the EV. While this normalization with respect to the EV is a natural choice when planning for the EV, it does not lead to viewpoint-invariant predictions for SVs [131]. In addition, it complicates the modeling of interactions among SVs. For instance, this coordinate frame makes it easy to infer which vehicles are close to the EV or in front of it (small magnitude of $x$ position). However, finding SVs that are close to each other is a second-order relation that is harder to learn. This is why other models [132, 133, 75, 103] process a scene from the viewpoint of each agent. Thus, the processing is viewpoint-invariant and makes it easier to represent interactions between all agents. On the downside, these models cannot use a shared representation for prediction and planning, and computational complexity scales linearly with the number of agents [6]. Moreover, per-agent coordinate frames are unsuitable for joint prediction, as each agent is processed individually.